OpenStudy (dsena):

How do you find the LCD of 3x and 9x^2?

OpenStudy (mathmale):

Are you in a situation where you have two fractions, one of which has denominator 3x and the other 9x^2? A B ------ - ------- 3x 9x^2 If so, you need to determine which factors 3x and 9x^2 have in common. Can 9x^2 be divided by 3x? Yes. So, your denominators are 3x and 3x(3x). The 2nd one, 3x3x, or 9x^2, is your LCD. Mult. A and 3x by 3x to obtain a fraction with the LCD in it that is equivalent to A/(3x).

OpenStudy (dsena):

Okay, thanks! The second denominator was 9x^2y actually. Would the LCD be 9x^2y?

OpenStudy (mathmale):

Yes, it would. If you want further feedback, please share your work here.

OpenStudy (dsena):

Okay, thank you for the help!

OpenStudy (mathmale):

You're welcome!

OpenStudy (dsena):

So I'm being asked to fill in the blank to make an equivalent fraction with the given denominator. \[\frac{ 5 }{ 3x}=\frac{ ? }{ 9x^2y }\]

OpenStudy (dsena):

Would I multiply the 9x^2y by 5?

OpenStudy (danjs):

you want the fraction on the right to be the same as the one on the left...

OpenStudy (danjs):

you see the denominator was multiplied by 3*x*y to get to the right fraction. To keep it equals the same, you have to do the same to the numerator

OpenStudy (danjs):

\[\large \frac{ 5 }{ 3x }*\frac{ 3xy }{ 3xy }=\]